# A Survey of Privacy-Preserving Model Explanations: Privacy Risks, Attacks, and Countermeasures

### v1

@@ 6.3.4. Attack success
Balanced accuracy (BA). This metric measures the accuracy of attack (e.g. membership inference), on a balanced dataset of members and non-members (Pawelczyk et al., 2023; Liu et al., 2024b):
@@
TPR at Low FPR. TPR at Low FPR (Liu et al., 2024b; Huang et al., 2023) is used to measure attack performance at a fixed FPR (e.g., 0.1%). Evaluating the True Positive Rate (TPR) at low False Positive Rates (FPR) is essential in scenarios where the cost of false positives is high, because it ensures that the positive results are both accurate and reliable. Low FPR evaluation is crucial particularly in imbalanced datasets, where false positives can outnumber true positives. For example, if a membership inference attack can pinpoint even a minuscule fraction of the training dataset with high precision, then the attack ought to be deemed effective (Pawelczyk et al., 2023).
@@ 7. Future Research Directions
7.1. Ethical Impacts
The push for explanations of automated models has spurred the creation of specialised tools and startups such as MS InterpretML, Fiddler Explainable AI Engine, IBM Explainability 360, Facebook Captum AI, and H2O Driverless AI. (Gade et al., 2019). Our survey delves into the unintended negative consequences of making ML models explainable, particularly highlighting the privacy risks associated with model explanations. Although the research is in its early stages, the threat is real. One immediate concern is the possibility that malicious entities could exploit existing findings as a basis for attacks on operational ML systems. Should such tools be implemented for high-risk data, like medical records or financial transactions, a thorough examination of their privacy implications becomes crucial. Another equally significant concern is the possible inherent conflict between the right to explain ML models (Goodman and Flaxman, 2017) and the preservation of user privacy. This dilemma warrants broader discussion beyond the ML research community, involving legal experts and policymakers (Banisar, 2011). Additionally, our findings suggest that the tension between explainability and privacy might disproportionately affect minority groups, either by exposing their data or providing them with lower-quality explanations (Shokri et al., 2021).
@@ 7.2. Privacy Tradeoffs
7.3. Unexplored Privacy Attacks
@@
7.4. Unexplored Model Explanations
Gillenwater et al. (Gillenwater et al., 2021) introduces a novel method for computing multiple quantiles in sensitive data with differential privacy. Traditional methods compromise on accuracy by either splitting the privacy budget across quantiles or inefficiently summarizing the entire distribution. The proposed approach uses an exponential mechanism to estimate multiple quantiles efficiently, achieving better accuracy and efficiency compared to existing methods. This is particularly relevant because there are emerging explainability measures based on quantiles (Ghosh et al., 2022; Li and van Leeuwen, 2023; Merz et al., 2022). Alvarez et al. (Alvarez Melis and Jaakkola, 2018) proposes the concept of self-explaining models that incorporate interpretability from the onset of learning. The authors design self-explaining models in a stepwise manner, starting from simple linear classifiers and advancing to more complex structures with built-in interpretability (Zhang et al., 2022). They introduce specialized regularization techniques to maintain faithfulness and stability. Olatunji et al. (Olatunji et al., 2023) pioneer the examination of privacy risks tied to feature explanations in graph neural networks (GNNs), presenting scenarios where adversaries attempt to unveil hidden relationships within the data, despite having limited access to the network’s structure (Khosla, 2022). The paper delves into various explanation methods for GNNs such as gradient-based, perturbation-based, and surrogate methods. Furthermore, it outlines potential adversarial attacks aimed at exploiting these explanations to compromise privacy and introduces a novel defense mechanism based on perturbing explanation bits to adhere to differential privacy standards. Other works (Tiddi and Schlobach, 2022; Rajabi and Etminani, 2022) examine the role of knowledge graphs as model explanations, positing that integrating structured, domain-specific knowledge can lead to more understandable, insightful, and trustworthy AI systems (Hung et al., 2019; Zhao et al., 2021a). However, knowledge graphs can be used to fuel privacy attacks such as de-anonymisation and membership inference (Qian et al., 2017; Wang et al., 2021).
7.5. Unexplored Data Modalities
Graph Data. The rapid development in the area of graph neural networks (GNNs) highlights a special treatment for GNN explainability (Wu et al., 2020; Huynh et al., 2021; Duong et al., 2022; Nguyen et al., 2015b, 2014). Yuan et al. (Yuan et al., 2022) discusses explainability methods specifically designed for Graph Neural Networks (GNNs) such as gradients/features-based, perturbation-based, surrogate, and decomposition methods. Prado et al. (Prado-Romero et al., 2023) provides a comprehensive overview of graph counterfactual explanations for GNNs. Privacy attacks on GNNs is also an emerging direction (Dai et al., 2022).
@@ 7.6. Privacy-Preserving Models
Exploring how privacy-preserving models, such as differentially private decision trees, reduce the success of privacy attacks represents a valuable research direction (Ferry et al., 2023a; Liu et al., 2022a, 2023a). Li et al. (Li et al., 2023) presents an Adaptive Differential Privacy (ADP) mechanism aimed at improving the interpretability of machine learning models without compromising privacy. This mechanism selectively injects noise into the less critical weights of a model’s parameters, thereby preserving the interpretability of important features which conventional differential privacy methods may obscure.
@@ 8. Conclusions
As the prevalence of model explanations grows, there is an emerging interest in understanding its repercussions, including aspects of fidelity, fairness, stability, and privacy. This survey offers a thorough investigation into the latest privacy-centric attacks on model explanations, establishing a comprehensive classification of these attacks based on their traits. Furthermore, it delves deeply into the present advanced research on defensive strategies and privacy-focused model explanations, uncovering common privacy design approaches and their variations.

### v2

@@ 6.3.4. Attack success
Balanced accuracy (BA). This metric measures the accuracy of attack (e.g. membership inference), on a balanced dataset of members and non-members (Pawelczyk et al., 2023; Liu et al., 2024d):
@@
TPR at Low FPR. TPR at Low FPR (Liu et al., 2024d; Huang et al., 2023) is used to measure attack performance at a fixed FPR (e.g., 0.1%). Evaluating the True Positive Rate (TPR) at low False Positive Rates (FPR) is essential in scenarios where the cost of false positives is high, because it ensures that the positive results are both accurate and reliable. Low FPR evaluation is crucial particularly in imbalanced datasets, where false positives can outnumber true positives. For example, if a membership inference attack can pinpoint even a minuscule fraction of the training dataset with high precision, then the attack ought to be deemed effective (Pawelczyk et al., 2023).
@@ 7. Future Research Directions
7.1. Ethical Implications
The push for explainable AI has led to the development of tools and startups like MS InterpretML, Fiddler Explainable AI Engine, IBM Explainability 360, Facebook Captum AI, and H2O Driverless AI (Gade et al., 2019). Our survey explores the privacy risks of making ML models explainable, highlighting the potential for malicious exploitation of these explanations, especially for high-risk data such as medical records and financial transactions. This raises concerns about the conflict between the right to explain ML models (Goodman and Flaxman, 2017) and user privacy, necessitating discussions involving legal experts and policymakers (Banisar, 2011). Additionally, the tension between explainability and privacy may disproportionately impact minority groups by either exposing their data or providing lower-quality explanations (Shokri et al., 2021).
@@ 7.1. Ethical Implications
Looking into the future, the ethical implications of privacy-preserving techniques include balancing privacy protection with transparency and fairness (Hu et al., 2022a). Techniques like differential privacy and federated learning secure data by adding noise or decentralising processing, but they can reduce model accuracy and transparency, complicating trust and understanding (Liu et al., 2024c, b). These methods can also introduce biases, affecting certain groups disproportionately and amplifying discrimination (Mi et al., 2024). Ensuring informed consent and user autonomy is crucial, necessitating clear communication about how these techniques impact data use and model performance (Zhang et al., 2024).
7.2. Regulatory Compliance
Privacy attacks on model explanations pose significant challenges under regulatory frameworks like the GDPR, which emphasise the protection of personal data and transparency in automated decision-making. Such attacks can lead to unauthorised data disclosure, complicating compliance with GDPR’s requirements for data subject rights, including access and erasure (Nguyen et al., 2022; Huynh et al., 2024). Additionally, privacy-preserving techniques that obscure model explanations may hinder transparency, making it difficult for organisations to demonstrate compliance and for individuals to understand AI decisions, thereby affecting accountability (Liu et al., 2024a). Moreover, these techniques must balance privacy and utility, as overly restrictive measures can impact the effectiveness and fairness of AI systems, posing further challenges for legal and ethical standards (Zhang et al., 2024).
@@ 7.3. Privacy Tradeoffs
7.4. Underexplored Privacy Attacks
@@
7.5. Underexplored Model Explanations
Gillenwater et al. (Gillenwater et al., 2021) introduces a novel method for computing multiple quantiles in sensitive data with differential privacy. Traditional methods compromise on accuracy by either splitting the privacy budget across quantiles or inefficiently summarizing the entire distribution. The proposed approach uses an exponential mechanism to estimate multiple quantiles efficiently, achieving better accuracy and efficiency compared to existing methods. This is particularly relevant because there are emerging explainability measures based on quantiles (Ghosh et al., 2022; Li and van Leeuwen, 2023; Merz et al., 2022). Alvarez et al. (Alvarez Melis and Jaakkola, 2018) proposes the concept of self-explaining models that incorporate interpretability from the onset of learning. The authors design self-explaining models in a stepwise manner, starting from simple linear classifiers and advancing to more complex structures with built-in interpretability (Zhang et al., 2022). They introduce specialized regularization techniques to maintain faithfulness and stability. Olatunji et al. (Olatunji et al., 2023) pioneer the examination of privacy risks tied to feature explanations in graph neural networks (GNNs), presenting scenarios where adversaries attempt to unveil hidden relationships within the data, despite having limited access to the network’s structure (Khosla, 2022). The paper delves into various explanation methods for GNNs such as gradient-based, perturbation-based, and surrogate methods. Furthermore, it outlines potential adversarial attacks aimed at exploiting these explanations to compromise privacy and introduces a novel defense mechanism based on perturbing explanation bits to adhere to differential privacy standards. Other works (Tiddi and Schlobach, 2022; Rajabi and Etminani, 2022) examine the role of knowledge graphs as model explanations, positing that integrating structured, domain-specific knowledge can lead to more understandable, insightful, and trustworthy AI systems. However, knowledge graphs can be used to fuel privacy attacks such as de-anonymisation and membership inference (Qian et al., 2017; Wang et al., 2021).
7.6. Underexplored Data Modalities
Graph Data. The rapid development in the area of graph neural networks (GNNs) (Huynh et al., 2021; Duong et al., 2022; Nguyen et al., 2014, 2015b; Hung et al., 2019) highlights a special treatment for GNN explainability (Wu et al., 2020). Yuan et al. (Yuan et al., 2022) discuss explainability methods specifically designed for Graph Neural Networks (GNNs) such as gradients/features-based, perturbation-based, surrogate, and decomposition methods. Prado et al. (Prado-Romero et al., 2023) provides a comprehensive overview of graph counterfactual explanations for GNNs. Privacy attacks on GNNs are also an emerging direction (Dai et al., 2022).
@@ 7.7. Privacy-Preserving Models
Exploring how privacy-preserving models, such as differentially private decision trees, reduce the success of privacy attacks represents a valuable research direction (Ferry et al., 2023a). Li et al. (Li et al., 2023) presents an Adaptive Differential Privacy (ADP) mechanism aimed at improving the interpretability of machine learning models without compromising privacy. This mechanism selectively injects noise into the less critical weights of a model’s parameters, thereby preserving the interpretability of important features which conventional differential privacy methods may obscure.
@@ 7.8. Privacy-Protecting Explanations
7.9. Time Complexity
Time complexity is crucial in privacy attacks on model explanations. Fast run-time methods pose higher risks by enabling rapid exploitation, while more complex iterative attacks are less practical due to longer execution times. The feasibility of these attacks depends on computational resources and scalability. Effective countermeasures must balance protection and performance to mitigate risks from fast, real-time attacks. Unfortunately, only a few works thoroughly discuss time complexity such as Shapley approximation (Jia et al., 2019a) and DP-quantiles (Gillenwater et al., 2021).
@@ 8. Conclusions
Summary. As the prevalence of model explanations grows, there is an emerging interest in understanding its repercussions, including aspects of fidelity, fairness, stability, and privacy. This survey offers a thorough investigation into the latest privacy-centric attacks on model explanations, establishing a comprehensive classification of these attacks based on their traits. Furthermore, it delves deeply into the present advanced research on defensive strategies and privacy-focused model explanations, uncovering common privacy design approaches and their variations.
@@
Challenges. The challenges for new work in this field, as highlighted in the survey, include:
  • Balancing Transparency and Privacy: Providing detailed explanations improves transparency but increases the risk of privacy breaches by revealing sensitive information embedded in the training data.
  • Granularity of Explanations: Detailed explanations can lead to direct inferences about data points, making it challenging to protect privacy without losing interpretability.
  • Understanding Privacy Leaks: Identifying the causes of privacy leaks through model explanations is complex and requires thorough investigation of different explanation methods and their vulnerabilities.
  • Diverse Attack Models: Developing comprehensive defenses against a wide range of privacy attacks, including membership inference, model inversion, and reconstruction attacks, is necessary but challenging due to the evolving nature of these attacks.
  • Countermeasure Effectiveness: Evaluating and improving the effectiveness of countermeasures, such as differential privacy and perturbation techniques, to ensure they do not compromise the utility of model explanations.
  • Dynamic Interaction Scenarios: Assessing the impact of repeated interactions between adversaries and the model in dynamic settings adds complexity to designing robust privacy-preserving methods.
  • Interpretable Surrogates: Surrogate models used for providing explanations can themselves become targets for privacy attacks, necessitating additional safeguards.
  • Scalability and Practicality: Implementing privacy-preserving techniques in real-world must balance scalability and practicality without significantly affecting model performance.